I restarted the LSC models in the usual way via the c1lsc reboot script. After doing this I was able to lock the YARM configuration for more noise coupling scripting.
The PMC and IMC were unlocked. Both were re-locked, and alignment of both cavities were adjusted so as to maximize MC2 trans (by hand, input alignment to PMC tweaked on PSL table, IMC alignment tweaked using slow bias voltages). I disabled the inputs to the WFS loops, as it looks like they are not able to deal with the glitching IMC suspensions. c1lsc models have crashed again but I am not worrying about that for now.
9pm: The alignment is wandering all over the place so I'm just closing the PSL shutter for now.